# We can catalyse it for you thermally: Catalytic transformations for thermal operations

### v1

@@ Abstract
What are the fundamental limits and advantages of using a catalyst to aid thermodynamic transformations between quantum systems? In this work, we answer this question by focusing on transformations between energy-incoherent states under the most general energy-conserving interactions among the system, the catalyst, and a thermal environment. The sole constraint is that the catalyst must return unperturbed and uncorrelated with the other subsystems. More precisely, we first characterise the set of states to which a given initial state can thermodynamically evolve (the catalysable future) or from which it can evolve (the catalysable past) with the help of a strict catalyst. Secondly, we derive lower bounds on the dimensionality required for the existence of catalysts under thermal process, along with bounds on the catalyst’s state preparation. Finally, we quantify the catalytic advantage in terms of the volume of the catalysable future and demonstrate its utility in an exemplary task of generating entanglement using thermal resources.
@@ I Introduction
Under this paradigm, the existence of a family of generalised free energies, which indicate what states can be reached from a given one (referred to as the present state), naturally decomposes the space of states into three distinct parts: the set of states to which the present state can evolve, known as the future thermal cone; the set of states from which the present state can be reached, or the past thermal cone; and the set of states that are neither in the past nor the future thermal cone, constituting the incomparable region [5, 6] – three regions analogous to the structure of the light cone in special relativity, which divides the space–time into future, past, and space-like region (see Fig. 1). Consequently, the issue of thermodynamic transformations in finite-size systems is characterised by the emergence of a finite-size effect known as incomparability.
@@
Arguably, one of the fundamental problems in quantum thermodynamics concerns quantification of the catalytic advantages in a given transformation. This issue can be approached by investigating the behaviour of the thermal cones when a catalyst is introduced. In this scenario, the catalyst cannot change future into the past or vice-verse, nor can it shrink both regions. Consequently, the growth of the past and future thermal cones occurs at the expense of the incomparable region. However, determining the set of states that can be achieved with the aid of a catalyst has been considered a highly non-trivial problem. The first steps for a related problem were made in Ref. [46, 47] within the realm of entanglement theory, where the authors focused on pure state transformations governed by entanglement-assisted local operations and classical communication (ELOCC). Therein, the authors consider bounds for the entanglement of a catalyst necessary to catalyse transformation between a given pair of otherwise incomparable states. Furthermore, they established a lower bound for the dimension of the catalyst required for a specific ELOCC transformation. As these results rely on the notion of majorisation, they automatically extend to the resource theory of coherence, as well as to so-called noisy operations, or incoherent thermal operations at infinite temperature.
@@
We answer this question by providing for any given starting state $\boldsymbol{p}$ a set of states $\mathcal{C}_{+}(\boldsymbol{p})$, called the catalysable future region of state $\boldsymbol{p}$, into which it can be transformed with the aid of a catalyst. The definition of catalysable past region $\mathcal{C}_{-}(\boldsymbol{p})$ follows analogously by taking $\boldsymbol{p}$ to be the target of a transformation to be catalysed. More specifically, for a $d$-dimensional energy-incoherent state, we provide an explicit construction of its catalysable past and future, along with a characterisation of the extreme points of the catalysable future. These results reveal fundamental limits on the advantages achievable with uncorrelated catalysts [see Fig. 2 for a warm-up example, where we present the future and past regions with and without catalyst for a three-and four-level systems]. As these results are general and do not rely on any specific assumptions regarding the catalytic state, our second focus is on the dimensionality and populations of the catalyst. We derive lower bounds on the dimensionality required for catalysts under thermal operations, together with bounds on the catalysts population. Our results shed light on how to construct a catalyst given two incomparable states. Surprisingly, our findings show that the no-go result from [47] formulated for the resource theory of entanglement and applicable for the infinite-temperature setting, stating that catalysis has no effect under local operations and classical communication for a main system of dimension three, ceases to apply in the context of thermodynamics at finite temperatures.
@@
The paper is organised as follows. In Section II, we introduce the resource-theoretic approach to thermodynamics, review known results concerning the conditions for state transformation under thermal operations, and establish the notation of the paper. Section III, collecting our main results, begins with the construction of the catalysable past and future regions, giving a simple method for generating all the extreme points of the catalysable future, thus providing the closest analogue to the Birkhoff theorem for the problem of catalysis in thermal operations. Next, we state our second main result, which establishes a lower bound on the dimensionality of the catalyst and introduces a constraint on its population vector for any pair of incomparable states. In Section IV we discuss a quantification of the catalytic advantage by volume of the catalysable future and past regions. Then, in Section V, we provide an application of our results to entanglement generation under thermal operations. Finally, we conclude with an outlook in Section VI. The technical derivation of our results can be found in the Appendices.
@@ II.1 Thermal operations
where $U$ is a joint unitary that commutes with the total Hamiltonian of the system and the bath $[U,H\otimes\mathbbm{1}_{E}+\mathbbm{1}\otimes H_{E}]=0$.
The fundamental question within the resource-theoretic approach is to identify the set of states that a given state $\rho$ can be transformed to under thermal operations. The reachability of states under TOs can be studied by introducing the notion of thermal cones [6]. Given $\rho$, the set of states achievable under a thermal operation is referred to as the future thermal cone ${\cal T}^{,\beta}_{+}(\rho)$. Conversely, the set of states that can evolve into $\rho$ is called the past thermal cone ${\cal T}^{,\beta}_{-}(\rho)$. The set of states that are neither in the past nor the future of $\rho$ is termed the incomparable thermal region ${\cal T}^{,\beta}_{\emptyset}(\rho)$. The general characterisation of thermal cones is not known beyond the simplest qubit case [53, 54].
@@
where $F_{\alpha}(\rho):=[D_{\alpha}(\rho\|\sigma)-\log Z]/\beta$, with $D_{\alpha}(\rho\|\sigma)$ being the quantum Rényi divergence [56, 57]. Nevertheless, a complete characterisation of the catalytic future thermal cone has not yet been addressed.
@@ II.2 Mathematical preliminaries
Consequently, the concept of $\beta$-ordering simplifies to arranging the probability vectors in non-increasing order. Thus, thermomajorisation reduces to the well-known concept of majorisation [58]. Second, thermomajorisation does not introduce a total order, a given pair of states $\boldsymbol{p}$ and $\boldsymbol{q}$ is said to be incomparable when neither $\boldsymbol{p}$ thermomajorises $\boldsymbol{q}$, nor $\boldsymbol{q}$ thermomajorises $\boldsymbol{p}$. We denote incomparable states as $\boldsymbol{p}\perp^{\beta}\boldsymbol{q}$.
@@ III.1 Catalysable past and future regions
We start by defining the catalysable sets:
Definition 5 (Catalysable region).
For a $d$-dimensional energy-incoherent state $\boldsymbol{p}$, we define the catalysable future (past) region $\mathcal{C}_{\pm}(\boldsymbol{p})$ as subsets of the incomparable region $\mathcal{T}_{\emptyset}(\boldsymbol{p})$ such that
$\mathcal{C}_{+}(\boldsymbol{p})$$\equiv\quantity{\boldsymbol{q}\in\Delta_{d}:\boldsymbol{q}\perp_{\beta}\boldsymbol{p},\,\exists_{k\geq 2,\,\boldsymbol{c}\in\Delta_{k}}:\boldsymbol{p}\otimes\boldsymbol{c}\succ_{\beta}\boldsymbol{q}\otimes\boldsymbol{c}},$ (16)
$\mathcal{C}_{-}(\boldsymbol{p})$$\equiv\quantity{\boldsymbol{q}\in\Delta_{d}:\boldsymbol{q}\perp_{\beta}\boldsymbol{p},\,\exists_{k\geq 2,\,\boldsymbol{c}\in\Delta_{k}}:\boldsymbol{q}\otimes\boldsymbol{c}\succ_{\beta}\boldsymbol{p}\otimes\boldsymbol{c}}.$ (17)
In what follows, if $\boldsymbol{q}\in\mathcal{C}_{+}(\boldsymbol{p})$, we will say that transformation $\boldsymbol{p}\rightarrow\boldsymbol{q}$ is catalysable, or that $\boldsymbol{q}$ is catalysable, omitting explicit relation to $\boldsymbol{p}$ whenever it is clear from the context.
The following result establishes a necessary condition for any state $\boldsymbol{q}$ to belong to the catalysable future region of a state $\boldsymbol{p}$:
Lemma 1 (Catalysability condition).
Consider a pair of incomparable states $\boldsymbol{q}\perp_{\beta}\boldsymbol{p}$. The state $\boldsymbol{q}$ is catalysable (with respect to $\boldsymbol{p}$), ie. $\boldsymbol{q}\in\mathcal{C}_{+}(\boldsymbol{p})$ [or, equivalently, $\boldsymbol{p}\in\mathcal{C}_{-}(\boldsymbol{q})$], only if it satisfies:
[…]
@@
As a result, the catalysable past and future can be explicitly constructed by using Lemma 1 together with the tangent vectors introduced in Sec. II.2.
Theorem 2 (Catalysable region characterisation).
@@
where $\mathcal{S}_{d}$ is the set of all possible $d$-element permutations. The catalysable future and past regions can then be reformulated as
@@
Finally, the extremal points of the above set can be derived directly from its thermomajorisation curve:
Corollary 3 (Extreme points of the catalysable future).
@@
It is interesting to note that the full catalysable future of any given state will have much fewer extreme points than the corresponding future thermal cone.
Lemma 4.
@@ III.2 Dimensionality bounds for thermal catalyst
The full proof of the above theorem is given in Appendix B. Therein, we present a heuristic proof in which we treat (thermo)majorisation curves as continuous objects, and an exact proof based on the notion of an embedding map [52], where thermomajorisation $\succ_{\beta}$ is equivalent to majorisation $\succ$ in the (embedded) higher-dimensional space. Both approaches agree in the obtained extension of the results presented in [47].
@@
$1-q_{1}+q_{1}\frac{\gamma_{1}-\gamma_{2}}{\gamma_{1}}$ $\leq 1-p_{2}$ (36)
@@
Figure 5: Isovolumetric curves for $d=3$. For a three-level system described by an equidistant energy spectrum with $E_{1}=0$, $E_{2}=1$, and $E_{3}=2$, we plot the volume of the catalysable future for several regimes of $\beta$. The lines indicate states with equal volumes, and the colors represent their magnitudes as a fraction of the maximal volume, $\mathcal{V}_{+}/\quantity(\max_{\Delta_{3}}\mathcal{V}_{+})$. The yellow star depicts the thermal state, and the dashed line indicates the boundaries between the six possible $\beta$-orderings, dividing the simplex into six chambers. Note that at the boundaries $\mathcal{V}_{+}$ goes to zero, while maxima are located at the outer edges of each chamber.
IV Quantifying catalytic advantages
The characterisation of the catalysable regions naturally enables the quantification of the catalytic advantage through their respective relative volumes. Given a state $\boldsymbol{p}$, we define

### v2

# Catalytic transformations for thermal operations

@@ Abstract
What are the fundamental limits and advantages of using a catalyst to aid thermodynamic transformations between quantum systems? In this work, we answer this question by focusing on transformations between energy-incoherent states under the most general energy-conserving interactions among the system, the catalyst, and a thermal environment. The sole constraint is that the catalyst must return unperturbed and uncorrelated with the other subsystems. More precisely, we first upper bound the set of states to which a given initial state can thermodynamically evolve (the catalysable future) or from which it can evolve (the catalysable past) with the help of a strict catalyst. Secondly, we derive lower bounds on the dimensionality required for the existence of catalysts under thermal process, along with bounds on the catalyst’s state preparation. Finally, we quantify the catalytic advantage in terms of the volume of the catalysable future and demonstrate its utility in an exemplary task of generating entanglement and cooling a quantum system using thermal resources.
@@ I Introduction
Under this paradigm, the existence of a family of generalised free energies, which indicate what states can be reached from a given one (referred to as the present state), naturally decomposes the space of states into three distinct parts: the set of states to which the present state can evolve, known as the future thermal cone $\mathcal{T}_{+}$; the set of states from which the present state can be reached, or the past thermal cone $\mathcal{T}_{-}$; and the set of states that are neither in the past nor the future thermal cone, constituting the incomparable region $\mathcal{T}_{\emptyset}$ [5, 6] – three regions analogous to the structure of the light cone in special relativity, which divides the space–time into future, past, and space-like region (see Fig. 1). Consequently, the issue of thermodynamic transformations in finite-size systems is characterised by the emergence of a finite-size effect known as incomparability.
@@
Arguably, one of the fundamental problems in quantum thermodynamics concerns quantification of the catalytic advantages in a given transformation. This issue can be approached by investigating the behaviour of the thermal cones when a catalyst is introduced. In this scenario, the catalyst cannot change future into the past or vice-verse, nor can it shrink both regions. Consequently, the growth of the past and future thermal cones occurs at the expense of the incomparable region. However, determining the set of states that can be achieved with the aid of a catalyst has been considered a highly non-trivial problem.
A complementary problem is that almost nothing is known about the precise form of the catalyst state required for a given transformation. Although in many cases the necessary (and sometimes sufficient) conditions for the existence of a catalyst are known, the theory remains agnostic regarding the precise form and properties of the catalytic state itself. The first steps towards addressing this problem were made in Ref. [44, 45] within the realm of entanglement theory, where the authors focused on pure state transformations governed by entanglement-assisted local operations and classical communication (ELOCC). Therein, the authors consider bounds for the entanglement of a catalyst necessary to catalyse transformation between a given pair of otherwise incomparable states. Furthermore, they established a lower bound for the dimension of the catalyst required for a specific ELOCC transformation. As these results rely on the notion of majorisation, they automatically extend to the resource theory of coherence, as well as to so-called noisy operations, or incoherent thermal operations at infinite temperature. For finite temperatures, state transformations between specific quantum states (those that are diagonal in the energy eigenbasis) are described by a more general notion known as thermomajorisation order. This concept does not have an equivalent physical counterpart in either entanglement or coherence theories. Currently, understanding any relationship between the dimensionality or form of the catalyst in a given thermal process aided by a strict catalyst remains an open problem.
@@
This question can be tackled by defining the catalytic future thermal cone $\mathcal{T}_{\mathcal{C}+}$. This set represents the states to which a given initial state $\boldsymbol{p}$ can be transformed with the aid of a catalyst. Similarly, one can also define the catalytic past thermal cone $\mathcal{T}_{\mathcal{C}-}$ by considering $\boldsymbol{p}$ as the target state of a transformation to be catalysed. We shed light on the answer to this question by bounding the catalytic past and future cones – in other words, we construct a catalysable past and future. More precisely, for a $d$-dimensional energy-incoherent state, we present an explicit construction of the catalysable past $\mathcal{C}_{-}$ and the catalysable future $\mathcal{C}_{+}$, along with a characterization of its extreme points. These results reveal fundamental limits on the advantages achievable with uncorrelated catalysts [see Fig. 2 for a warm-up example, where we present the future and past regions with and without catalyst for a three-and four-level systems]. As these results are general and do not rely on any specific assumptions regarding the catalytic state, our second focus is on the dimensionality and populations of the catalyst. We derive lower bounds on the dimensionality required for catalysts under thermal operations, together with bounds on the catalysts population. Our results allows us to construct a catalytic state given two incomparable states. Surprisingly, our findings show that the no-go result from [45] formulated for the resource theory of entanglement and applicable for the infinite-temperature setting, stating that catalysis has no effect under local operations and classical communication for a main system of dimension three, ceases to apply in the context of thermodynamics at finite temperatures.
@@
The paper is organised as follows. In Section II, we introduce the resource-theoretic approach to thermodynamics, review known results concerning the conditions for state transformation under thermal operations, and establish the notation of the paper. Section III, collecting our main results, begins with the construction of the catalysable past and future regions, giving a simple method for generating all the extreme points of the catalysable future, thus providing the closest analogue to the Birkhoff theorem for the problem of catalysis in thermal operations. Next, we state our second main result, which establishes a lower bound on the dimensionality of the catalyst and introduces a constraint on its population vector for any pair of incomparable states. In Section III.3 we discuss a quantification of the catalytic advantage by volume of the catalysable future and past regions. Then, in Section IV.1, we provide an application of our results to entanglement generation under thermal operations. Then, in Section IV.2 we provide additional application in catalytic cooling. Finally, we conclude with an outlook in Section V. The technical derivation of our results can be found in the Appendices.
@@ II.1 Thermal operations
where $U$ is a joint unitary that commutes with the total Hamiltonian of the system and the bath $[U,H\otimes\mathbbm{1}_{\textsf{R}}+\mathbbm{1}\otimes H_{\textsf{R}}]=0$.
The fundamental question within the resource-theoretic approach is to identify the set of states that a given state $\rho$ can be transformed to under thermal operations. The reachability of states under TOs can be studied by introducing the notion of thermal cones [6]. Given $\rho$, the set of states achievable under a thermal operation is referred to as the future thermal cone ${\cal T}_{+}(\rho)$. Conversely, the set of states that can evolve into $\rho$ is called the past thermal cone ${\cal T}_{-}(\rho)$. The set of states that are neither in the past nor the future of $\rho$ is termed the incomparable thermal region ${\cal T}_{\emptyset}(\rho)$. The general characterization of thermal cones is not known beyond the simplest qubit case [51, 52].
@@
where $F_{\alpha}(\rho):=[D_{\alpha}(\rho\|\sigma)-\log Z]/\beta$, with $D_{\alpha}(\rho\|\sigma)$ being the quantum Rényi divergence [54, 55]. Nevertheless, a complete characterization of the catalytic future thermal cone has not yet been addressed.
@@ II.2 Mathematical preliminaries
Consequently, the concept of $\beta$-ordering simplifies to arranging the probability vectors in non-increasing order. Thus, thermomajorisation reduces to the well-known concept of majorisation [56]. Second, thermomajorisation does not introduce a total order (or even fails to be a partial order in some cases, as proven in Ref. [57]), a given pair of states $\boldsymbol{p}$ and $\boldsymbol{q}$ is said to be incomparable when neither $\boldsymbol{p}$ thermomajorises $\boldsymbol{q}$, nor $\boldsymbol{q}$ thermomajorises $\boldsymbol{p}$. We denote incomparable states as $\boldsymbol{p}\perp^{\beta}\boldsymbol{q}$.
@@ III.1 Catalysable past and future regions
We start by defining the catalytic thermal cones:
Definition 5 (Catalytic thermal cones).
For a $d$-dimensional energy-incoherent state $\boldsymbol{p}$, we define the catalytic future/past thermal cone $\mathcal{T}_{\mathcal{C}{+/-}}$ as subsets of the incomparable region $\mathcal{T}_{\emptyset}(\boldsymbol{p})$ such that
$\mathcal{T}_{\mathcal{C}+}(\boldsymbol{p})$$\equiv\quantity{\boldsymbol{q}\in\Delta_{d}:\boldsymbol{q}\perp_{\beta}\boldsymbol{p},\exists_{k\geq 2,\boldsymbol{c}\in\Delta_{k}}:\boldsymbol{p}\otimes\boldsymbol{c}\succ_{\beta}\boldsymbol{q}\otimes\boldsymbol{c}},$ (16)
$\mathcal{T}_{\mathcal{C}-}(\boldsymbol{p})$$\equiv\quantity{\boldsymbol{q}\in\Delta_{d}:\boldsymbol{q}\perp_{\beta}\boldsymbol{p},\exists_{k\geq 2,\boldsymbol{c}\in\Delta_{k}}:\boldsymbol{q}\otimes\boldsymbol{c}\succ_{\beta}\boldsymbol{p}\otimes\boldsymbol{c}},$ (17)
In what follows, if $\boldsymbol{q}\in\mathcal{T}_{\emptyset}(\boldsymbol{p})$, we will say that transformation $\boldsymbol{p}\rightarrow\boldsymbol{q}$ is catalysed, or that $\boldsymbol{q}$ is catalysed, omitting explicit relation to $\boldsymbol{p}$ whenever it is clear from the context.
The following result establishes a necessary condition for any state $\boldsymbol{q}$ to belong to the catalytic future thermal cone of a state $\boldsymbol{p}$:
Lemma 1 (Catalytic condition).
Consider a pair of incomparable states $\boldsymbol{q}\perp_{\beta}\boldsymbol{p}$. The state $\boldsymbol{q}$ is catalysed (with respect to $\boldsymbol{p}$), ie. $\boldsymbol{q}\in\mathcal{T}_{\mathcal{C}+}(\boldsymbol{p})$ [or, equivalently, $\boldsymbol{p}\in\mathcal{T}_{\mathcal{C}-}(\boldsymbol{q})$], only if it satisfies:
$s_{1}\quantity(\boldsymbol{p})>s_{1}\quantity(\boldsymbol{q})\quad\text{and}\quad s_{d}\quantity(\boldsymbol{p})<s_{d}\quantity(\boldsymbol{q}).$ (18)
@@
Combining Lemma 1 with the tangent vectors introduced in Sec. II.2 enables us to provide a simple upper bound for the catalytic thermal cones and characterise the catalysable regions as:
Theorem 2 (Catalysable regions).
@@
where $\mathcal{S}_{d}$ is the set of all possible $d$-element permutations. The catalysable future and past regions are bounded by
@@
That is, $\mathcal{C}_{-/+}(\boldsymbol{p})\subset\mathcal{T}_{\mathcal{C}-/+}$.
@@
Theorem 2 establishes a fundamental limit for when an incoherent strict catalyst is allowed to be used in a thermodynamic process. In other words, there is no thermodynamic process aided by a strict catalyst that can bring the initial state out of the catalysable future $\mathcal{C}_{+}$. Hence, the advantages gained from using a catalyst are captured by such a region. However, it is important to emphasize the limitations of such a result, namely given a state $\boldsymbol{q}$ belonging to such a region, whether there exists a catalyst $\boldsymbol{c}$ and a thermal operation $\mathcal{E}$, such that $\mathcal{E}(\boldsymbol{p}\otimes\boldsymbol{c})=\boldsymbol{q}$ still remains an open problem.
@@
Finally, the extremal points of the catalysable set can be derived directly from its thermomajorisation curve:
Corollary 3 (Extreme points of $\mathcal{C}_{+}$).
@@
It is interesting to note that the full catalysable future $\mathcal{C}_{+}$ of any given state will have much fewer extreme points than the corresponding future thermal cone.
Lemma 4 (Number of extreme points).
@@
The concept of catalysable regions applies also to other majorisation-based resource theories, such as entanglement and coherence theories, where they proceed from the [45] directly in combination with the tangent vectors. These theories are defined by sets of free operations and free states: local operations and classical communication (LOCC) with separable states in entanglement theory [58], and incoherent operations (IO) with incoherent states in coherence theory [59]. In each theory, quantum states can be represented by probability distributions, which are crucial for defining state transformation conditions under free operations. However, in both cases, the partial order that emerges is precisely the opposite of the thermodynamic order in the infinite temperature limit [60, 61]. As a result, what constitutes the catalysable thermodynamic past and future in thermodynamics becomes the future and past in entanglement and coherence theories, while the incomparable region remains unchanged. Nevertheless, it is important to highlight that there is no fully equivalent physical situation in either entanglement or coherence theories where thermomajorisation is applicable.
@@ III.2 Dimensionality bounds for thermal catalyst
The full proof of the above theorem is given in Appendix B. Therein, we present a heuristic proof in which we treat (thermo)majorisation curves as continuous objects, and an exact proof based on the notion of an embedding map [50], where thermomajorisation $\succ_{\beta}$ is equivalent to majorisation $\succ$ in the (embedded) higher-dimensional space. Both approaches agree in the obtained extension of the results presented in [45]. Crucial insight leading to the results, however, is that catalysability of thermal states is encoded in the slopes of thermomajorisation functions, which may be obtained directly from the relation between the state of the system and the underlying Gibbs distribution. This stands in contrast with the standard understanding that transition from majorisation to thermomajorisation requires generically infinite-dimensional system.
@@
$1-q_{1}+q_{1}\frac{\gamma_{1}-\gamma_{2}}{\gamma_{1}}$ $\leq 1-p_{2},$ (36)
@@
Figure 5: Isovolumetric curves for $d=3$. For a three-level system described by an equidistant energy spectrum with $E_{1}=0$, $E_{2}=1$, and $E_{3}=2$, we plot the volume of the catalysable future for several regimes of $\beta$. The lines indicate states with equal volumes, and the colours represent their magnitudes as a fraction of the maximal volume, $\mathcal{V}_{+}/\quantity(\max_{\Delta_{3}}\mathcal{V}_{+})$. The yellow star depicts the thermal state, and the dashed line indicates the boundaries between the six possible $\beta$-orderings, dividing the simplex into six chambers. Note that at the boundaries $\mathcal{V}_{+}$ goes to zero, while maxima are located at the outer edges of each chamber.
III.3 Quantifying catalytic advantages
The characterization of the catalysable regions naturally enables the quantification of the catalytic advantage through their respective relative volumes. Given a state $\boldsymbol{p}$, we define